OpenStudy (anonymous):

determine whether the graphs are parrallel or perpendicular -y-7=-x and y-x=3

OpenStudy (anonymous):

look at the slop value = number in a front of x after you present Y in a terms of s: y=x-7 and y=x+3 the value/number in a front of x is 1, so the slope of the line is the same= lines are parallel